OpenStudy (anonymous):

4x=-3y+17 3x-4y=-6 What is the solutions this system of equations?

hero (hero):

Hint: Write out the system in this form: 4x + 3y = 17 3x - 4y = - 6 Combine both equations together Then isolate y

hero (hero):

(4x + 3x) + (3y - 4y) = (17 - 6)

hero (hero):

7x - y = 11 y = 7x - 11

hero (hero):

Substitute the expression for y back in to one of the original equations to solve for x.

OpenStudy (anonymous):

Is it 2,3

OpenStudy (anonymous):

Thanks